# Spike No More: Stabilizing the Pre-training of Large Language Models

### 28 Dec 2023 v1

@@ 4 Main Experiments
We verify empirical effectiveness of our theoretical analyses. In detail, we demonstrate that the method that prevents the exploding gradients also prevents loss and gradient spikes. In particular, we investigate the effectiveness of modifications to prevent the explosion due to LNs because the empirical effectiveness of the scaled initialization already has been revealed in previous studies (Shoeybi et al., 2020; Le Scao et al., 2022).
@@ 4.1 Datasets
We used C4 (Raffel et al., 2020) that consists of clean English texts extracted from Common Crawl [^7] https://commoncrawl.org/ as our LLM pre-training corpus. We also used the separated part of C4 as our validation data. We used GPT-2 vocabulary (Radford et al., 2019) that contains Byte Pair Encoding (BPE) subword units (Sennrich et al., 2016) as our vocabulary. For evaluation of each method, we computed perplexity on WikiText (Merity et al., 2017) and LAMBADA (Paperno et al., 2016) datasets.
@@ 4.2 Model Configurations
This configuration uses only the scaled initialization. Thus, this configuration prevents the explosion due to the residual branch but faces the explosion due to LNs.
@@
where $\gamma$ is a hyper-parameter and $\mathrm{Detach}$ detaches an input from the computational graph. We assign $0.1$ to $\gamma$ as in Zeng et al. (2023). Zeng et al. (2023) indicated that this method empirically prevent the loss spike. However, this method does not satisfy the requirements to prevent the explosion due to LNs, and thus we show that this method does not completely solve the loss spike.
@@ 4.3 Results
Figure 4 shows loss values of each method in validation data. Figure 5 shows gradient norms of each method. These figures indicate that Vanilla and Embed Detach faced loss and gradient spikes. In contrast, Embed LN and Scaled Embed did not face spikes. These results correspond to our theoretical analyses described in Section 3. Thus, we have to prevent the exploding gradients to avoid spikes in LLM pre-training.
In comparison between 350M and 1.7B parameters, spikes occurred more frequent in 1.7B parameters. Because we initialize embeddings with $\mathcal{N}(0,\sigma)$ where $\sigma=\sqrt{2/5d}$, the standard deviations of embeddings become small as $d$ gets larger. This means that the standard deviation of Equations (1) and (2) also become small as $d$ gets larger. As described in Section 3.2, the smaller the standard deviation the LN input is, the larger its gradient is. Therefore, if we increase $d$ without any technique to prevent the explosion due to LNs, a model become more unstable. This result corresponds to previous study report (Le Scao et al., 2022; Chowdhery et al., 2022; Zeng et al., 2023) that their model became more unstable as they increased the number of parameters.
@@
Table 1 shows perplexities of each method on WikiText and LAMBADA. This table shows that Embed LN and Scaled Embed achieved comparable performance. This result implies that methods have no significant difference from each other on their performance if each method prevents loss and gradient spikes. In contrast, perplexities of Vanilla and Embed Detach are worse except for Vanilla with 350M parameters in LAMBADA, and the difference on the performance is larger in a large amount of parameters. This result implies that addressing spikes has more serious influence on the performance as the parameter size gets larger. We discuss this matter more detail in Section 5.1.
@@ 5.1 Varying Learning Rate
Le Scao et al. (2022) reported that the stable method such as Embed LN was worse than Vanilla. However, in Section 4, the stable methods, Scaled Embed and Embed LN, achieved better performance than Vanilla in the 1.7B parameter configuration. We suppose that the difference of the learning rate causes this gap of findings. In this section, we tried to train Vanilla and Scaled Embed with larger and smaller learning rates: $\mathrm{lr}=1.0\times 10^{-3}$ and $1.0\times 10^{-4}$ respectively.
Figure 6 shows loss values of each configuration in validation data. As shown in this figure, the larger learning rate we used, the more frequent the spikes occurred in Vanilla. In particular, in $\mathrm{lr}=1.0\times 10^{-3}$, the training of Vanilla with 1.7B parameters failed because its gradient exploded as shown in Figure 7. In contrast, Scaled Embed stabilized the training, and thus, its loss values consistently decreased.
Table 2 shows perplexities of each configuration in evaluation data. This table indicates that Vanilla with 350M parameters achieved better performance in $\mathrm{lr}=1.0\times 10^{-4}$ that is the situation where its training did not face any spike. This result corresponds to the report of Le Scao et al. (2022). Thus, we suppose that they conducted the comparison in a too small learning rate to stabilize Vanilla. In contrast, the stable methods are more effective in training with a large learning rate as shown in Figure 6 and Table 2. Therefore, if Le Scao et al. (2022) used a relatively large learning rate in their experiments, their stable method achieved better performance.
@@ 5.2 Varying Sequence Length
As discussed in Section 3.2, in Vanilla, the smaller the standard deviation of each sub-layer is, the larger the gradient of its layer is. In addition, the standard deviation of the self-attention layer depends on the result of $\mathrm{softmax}$ function. We considered the maximum value case in Section 3.2. In contrast, in this section, we consider the minimum value case, i.e., the result of $\mathrm{softmax}$ is the uniform distribution. Let $L$ is the length of an input sequence. When the result of $\mathrm{softmax}$ is the uniform distribution, we multiply $1/L$ by each vector in $X$, and thus, their variances become $(1/L)^{2}$. Therefore, we obtain the minimum variance of $\mathrm{softmax}(\frac{xW_{Qi}(XW_{Ki})^{\mathrm{T}}}{\sqrt{d_{head}}})X$ with the following equation:
$\mathrm{min}(\mathrm{var}(\mathrm{softmax}(\frac{xW_{Qi}(XW_{Ki})^{\mathrm{T}}}{\sqrt{d_{head}}})X))$ $=\Bigl{(}\frac{1}{L}\Bigr{)}^{2}L,$
@@
We conducted experiments with varying $L$ from $128$ to $2048$. To use the same number of tokens to update parameters, we adjusted the batch size. Figure 8 shows loss values of Vanilla with each $L$ configuration in the validation data. This figure shows that spikes occurred only in the large $L$, i.e., $1024$ and $2048$. Moreover, the spikes are more probable to occur at the early stage of the pre-training. We suppose that the result of $\mathrm{softmax}$ in the early stage is similar to the uniform distribution because input vectors at the early stage are similar to random vectors. In other words, we suppose that the result of $\mathrm{softmax}$ at the early stage is difficult to be a peaked distribution because input vectors are not trained sufficiently. Therefore, using a short sequence stabilizes the training at the early stage as reported in Li et al. (2022).
@@ 5.3 RMSNorm
$||\frac{\delta\mathrm{RMSNorm}(x)}{\delta x}||_{2}=\mathcal{O}(\frac{\sqrt{d}}{||x||_{2}}).$ (16)
@@
Figures 9 and 10 show the loss values and gradient norms of the Vanilla configuration in Section 4 and the one using RMSNorms instead of LNs (“RMSNorm” in figures) with 350M parameters. We trained them with $\mathrm{lr}=5.0\times 10^{-4}$ as in Section 4888We tried to train them with $\mathrm{lr}=1.0\times 10^{-3}$ but RMSNorm exploded.. As shown in these figures, RMSNorm faced loss and gradient spikes in the similar location to ones of Vanilla. These empirical results also indicate that the RMSNorms have the same issue as LNs in terms of the exploding gradients.
@@ 6 Conclusion
This paper provides a reason why LLMs sometimes face loss spikes during their pre-training. Through theoretical analyses, we introduced two causes of the exploding gradients during LLM pre-training: the residual branch and LNs. We also described requirements to prevent the exploding gradients. To satisfy the requirements, we introduced the combination of simple modifications to embeddings and the scaled initialization that is widely used initialization method in LLM pre-training. Experiments indicated that the combination avoided loss spikes. Moreover, we can train the combined methods with a relatively larger learning rate, and thus, they can achieve better performance.
@@ Appendix B Related Work
To stabilize trainings of Transformer based neural language models, there have been various discussions on the architecture (Xiong et al., 2020; Liu et al., 2020; Takase et al., 2023; Zeng et al., 2023; Zhai et al., 2023), initialization method (Nguyen & Salazar, 2019; Zhang et al., 2019b; Huang et al., 2020; Wang et al., 2022), training strategy (Zhang et al., 2022; Li et al., 2022), and loss function (Chowdhery et al., 2022; Wortsman et al., 2023).
Xiong et al. (2020) theoretically analyzed gradient scales of each part in Transformers, and indicated that the Pre-LN Transformer is more stable than the Post-LN Transformer, that is the original Transformer architecture (Vaswani et al., 2017). Since the Pre-LN Transformer is more stable than the Post-LN Transformer theoretically and empirically, recent studies mainly have used the Pre-LN Transformer to construct a LLM. We also assume using the Pre-LN Transformer in the analysis on the training dynamics in this paper.
@@
For the initialization methods, Nguyen & Salazar (2019) proposed a strategy to initialize parameters of Transformers with small values to stabilize their trainings. Zhang et al. (2019b) and Huang et al. (2020) indicated that we can remove layer normalizations in Transformers if we used their proposed initialization methods. Wang et al. (2022) adjusted initial parameter scales based on the number of layers to stabilize the Post-LN Transformer. In this study, we indicated that an initialization method affects the LLM pre-training dynamics. In particular, we indicated that the widely used initialization method (Shoeybi et al., 2020) prevents the explosion due to the residual branch but causes the explosion due to layer normalizations. Moreover, we proved that we can prevent the explosion due to layer normalizations by making the variance of embeddings close to $1$.
@@
Rajbhandari et al. (2020) proposed ZeRO that reduces memory redundancies during the multi GPU training without increasing communication volume. Dao et al. (2022) focused on GPU memory reads/writes, and proposed FlashAttention that accelerates the speed of attention mechanisms in Transformers. To reduce the number of computation in the attention mechanism, Shazeer (2019) proposed the multi-query attention that shares one key and value across all of the attention heads in each layer. Takase & Kiyono (2023) explored several parameter sharing strategies, and indicated that parameter sharing across some layers can achieve comparable performance to the vanilla model with a small number of parameters. Moreover, several studies have explored a better construction way with a limited budget (Izsak et al., 2021; Takase & Kiyono, 2021). We believe that we can take advantage of their findings to make our LLMs more efficient.
@@ Appendix D Comparison with Post-LN Transformer
As described in Section 2.1, recent studies use the Pre-LN Transformer architecture to construct their LLMs because the architecture is more stable. In contrast, some recent studies reported that the Post-LN Transformer, which is the original architecture, can achieve better performance than the Pre-LN if we address the instability issue in the Post-LN, i.e., the vanishing gradient problem (Liu et al., 2020; Takase et al., 2023; Wang et al., 2022). We discuss whether the Pre-LN Transformer entirely underperforms the Post-LN. We conduct experiments on machine translation experiments because previous studies mainly focused on them.
We followed the experimental settings in Takase et al. (2023). Table 4 shows the details of hyper-parameters. We used the WMT English-to-German training dataset (Vaswani et al., 2017; Ott et al., 2018), and evaluated each model in newstest2010-2016. We used the encoder-decoder architecture proposed by Peitz et al. (2019). To stabilize the Post-LN Transformer, we applied DeepNet (Wang et al., 2022) and B2T connection (Takase et al., 2023). We compared them to Scaled Embed, that is the Pre-LN Transformer with the stabilizing techniques described in this paper.
Table 5 shows the averaged BLEU scores among newstest2010-2016. For the BLEU score calculation, we used SacreBLEU (Post, 2018) to obtain compatible scores (Marie et al., 2021). The signature of SacreBLEU is BLEU+case.mixed+numrefs.1+smooth.exp+tok.13a+version.1.5.0. As shown in this table, we used two learning rate: $\mathrm{lr}=1.0\times 10^{-3}$ and $3.0\times 10^{-3}$. For $\mathrm{lr}=1.0\times 10^{-3}$, DeepNet and B2T connection outperformed Scaled Embed. Thus, the Post-LN Transformer based methods achieved better performance than the Pre-LN Transformer based method. This result corresponds to reports in previous studies (Liu et al., 2020; Wang et al., 2022; Takase et al., 2023).
On the other hand, for $\mathrm{lr}=3.0\times 10^{-3}$, Scaled Embed achieved better performance than the others with $\mathrm{lr}=1.0\times 10^{-3}$, and the training of the others failed due to the exploding gradients. This result indicates that the Pre-LN Transformer based method can achieve better performance if we use a large learning rate. Therefore, the Pre-LN Transformer (with the stabilizing techniques) is more stable than the Post-LN Transformer based method, and thus, it can achieve better performance when we use a large learning rate that is too large to train the Post-LN Transformers.
@@ Appendix E Pre-training of the Model with 13B Parameters
We conducted pre-trainings of models with 13B parameters to indicate that our modification can stabilize a model with much more parameters than ones discussed in Section 4. To make this experiment close to a realistic situation, as shown in Table 3, we increased the batch size and the number of updates, and decreased the Adam $\beta_{2}$. In particular, for Adam $\beta_{2}$, most studies have used $0.95$ to stabilize their pre-trainings (Brown et al., 2020; Zhang et al., 2022; Zeng et al., 2023; Biderman et al., 2023; Touvron et al., 2023), and thus, we also used $0.95$ in this experiment. We tried two learning rates: $3.0\times 10^{-4}$, which is the same value in Touvron et al. (2023), and $1.0\times 10^{-4}$.
Figure 11 shows loss values of each configuration in validation data. As shown in (a) of this figure, the loss value of Vanilla rose from approximately 10000 steps in $\mathrm{lr}=3.0\times 10^{-4}$. Then, the gradient of this model became too large to continue its pre-training. In contrast, the loss value of Scaled Embed consistently decreased. This result indicates that Scaled Embed stabilized the pre-training. We emphasize that the pre-training of Vanilla is essentially unstable even if we use the widely used Adam $\beta_{2}$ value, $0.95$, which is known as the technique to stabilize the pre-training, and our modification is also effective for the stabilization in this realistic situation.
Table 6 shows perplexities of each configuration in evaluation data. This table indicates that we can achieve better performance when we use larger learning rate in the same as in Section 5.1. In addition, the perplexities of Scaled Embed were comparable to ones of Vanilla when we used the small learning rate: $\mathrm{lr}=1.0\times 10^{-4}$. These results imply that our modification does not have any considerable risk in pre-training. Thus, we recommend making the standard deviation of each embedding close to $1$ to stabilize the pre-trainings of LLMs.

### 02 Feb 2024 v2

@@ 4 Main Experiments
We verify the empirical effectiveness of our theoretical analyses. In detail, we demonstrate that the method that prevents the exploding gradients also prevents loss and gradient spikes. In particular, we investigate the effectiveness of modifications to prevent the explosion due to LNs because the empirical effectiveness of the scaled initialization has already been revealed in previous studies (Shoeybi et al., 2020; Le Scao et al., 2022).
@@ 4.1 Datasets
We used C4 (Raffel et al., 2020) that consists of clean English texts extracted from Common Crawl [^7] https://commoncrawl.org/ as our LLM pre-training corpus. We also used the separated part of C4 as our validation data. We used GPT-2 vocabulary (Radford et al., 2019) that contains Byte Pair Encoding (BPE) subword units (Sennrich et al., 2016) as our vocabulary. To evaluate each method, we computed perplexity on WikiText (Merity et al., 2017) and LAMBADA (Paperno et al., 2016) datasets.
@@ 4.2 Model Configurations
This configuration uses only the scaled initialization. Thus, this configuration prevents the explosion due to the shortcut but faces the explosion due to LNs.
@@
where $\gamma$ is a hyper-parameter and $\mathrm{Detach}$ detaches an input from the computational graph. We assign $0.1$ to $\gamma$ as in Zeng et al. (2023). Zeng et al. (2023) indicated that this method empirically prevents the loss spike. However, this method does not satisfy the requirements to prevent the explosion due to LNs, and thus, we show that this method does not completely solve the loss spike.
@@ 4.3 Results
Figure 4 shows the loss values of each method in validation data. Figure 5 shows the gradient norms of each method. These figures indicate that Vanilla and Embed Detach faced loss and gradient spikes. In contrast, Embed LN and Scaled Embed did not face spikes. These results correspond to our theoretical analyses described in Section 3. Thus, we must prevent the exploding gradients to avoid spikes in LLM pre-training.
In comparison between 350M and 1.7B parameters, spikes occurred more frequent in 1.7B parameters. Because we initialize embeddings with $\mathcal{N}(0,\sigma)$ where $\sigma=\sqrt{2/5d}$, the standard deviations of embeddings become small as $d$ gets larger. This means that the standard deviation of Equations (1) and (2) also become small as $d$ gets larger. As described in Section 3.2, the smaller the standard deviation the LN input is, the larger its gradient is. Therefore, if we increase $d$ without any technique to prevent the explosion due to LNs, a model becomes more unstable. This result corresponds to the previous study reports (Le Scao et al., 2022; Chowdhery et al., 2022; Zeng et al., 2023) that their model became more unstable as they increased the number of parameters.
@@
Table 1 shows the perplexities of each method on WikiText and LAMBADA. This table shows that Embed LN and Scaled Embed achieved comparable performance. This result implies that methods have no significant difference from each other in their performance if each method prevents loss and gradient spikes. In contrast, the perplexities of Vanilla and Embed Detach are worse except for Vanilla with 350M parameters in LAMBADA, and the difference in the performance is larger in a large amount of parameters. This result implies that addressing spikes has a more serious influence on the performance as the parameter size gets larger. We discuss this matter in more detail in Section 5.1.
@@ 5.1 Varying Learning Rate
Le Scao et al. (2022) reported that the stable method, such as Embed LN, was worse than Vanilla. However, in Section 4, the stable methods, Scaled Embed and Embed LN, achieved better performance than Vanilla in the 1.7B parameter configuration. We suppose that the difference in the learning rate causes this gap in findings. In this section, we tried to train Vanilla and Scaled Embed with larger and smaller learning rates: $\mathrm{lr}=1.0\times 10^{-3}$ and $1.0\times 10^{-4}$ respectively.
Figure 6 shows loss values of each configuration in validation data. As shown in this figure, the larger the learning rate we used, the more frequent the spikes occurred in Vanilla. In particular, in $\mathrm{lr}=1.0\times 10^{-3}$, the training of Vanilla with 1.7B parameters failed because its gradient exploded, as shown in Figure 7. In contrast, Scaled Embed stabilized the training, and thus, its loss values consistently decreased.
Table 2 shows the perplexities of each configuration in evaluation data. This table indicates that Vanilla with 350M parameters achieved better performance in $\mathrm{lr}=1.0\times 10^{-4}$ that is the situation where its training did not face any spike. This result corresponds to the report of Le Scao et al. (2022). Thus, we suppose that they conducted the comparison with a too-small learning rate to stabilize Vanilla. In contrast, the stable methods are more effective in training with a large learning rate as shown in Figure 6 and Table 2. Therefore, if Le Scao et al. (2022) used a relatively large learning rate in their experiments, their stable method achieved better performance.
@@ 5.2 Varying Sequence Length
As discussed in Section 3.2, in Vanilla, the smaller the standard deviation of each sub-layer is, the larger the gradient of its layer is. In addition, the standard deviation of the self-attention layer depends on the result of $\mathrm{softmax}$ function. We considered the maximum value case in Section 3.2. In contrast, in this section, we consider the minimum value case, i.e., the result of $\mathrm{softmax}$ is the uniform distribution. Let $L$ be the length of an input sequence. When the result of $\mathrm{softmax}$ is the uniform distribution, we multiply $1/L$ by each vector in $X$. Because $X$ is the pack of layer-normalized vectors, their variances become $(1/L)^{2}$ after the multiplication. Then, we compute the sum of the $L$ vectors whose variances are $(1/L)^{2}$. To obtain the variance after the summation, we multiply $L$ by the $(1/L)^{2}$ based on Equation (6). Formally, we obtain the minimum variance of $\mathrm{softmax}(Z)X$ with the following equation:
$\mathrm{min}(\mathrm{var}(\mathrm{softmax}(Z)X))$ $=\left(\frac{1}{L}\right)^{2}L$
@@
We conducted experiments with varying $L$ from $128$ to $2048$. To use the same number of tokens to update parameters, we adjusted the batch size. Figure 8 shows loss values of Vanilla with each $L$ configuration in the validation data. This figure shows that spikes occurred only in the large $L$, i.e., $1024$ and $2048$. Moreover, the spikes are more probable to occur at the early stage of the pre-training. We suppose that the result of $\mathrm{softmax}$ in the early stage is similar to the uniform distribution because input vectors at the early stage are similar to random vectors. In other words, we suppose that the result of $\mathrm{softmax}$ at the early stage is challenging to be a peaked distribution because input vectors are not trained sufficiently. Therefore, using a short sequence stabilizes the training at the early stage, as reported in Li et al. (2022).
@@ 5.3 RMSNorm
$\left\lVert\frac{\delta\mathrm{RMSNorm}(x)}{\delta x}\right\rVert_{2}=\mathcal{O}\left(\frac{\sqrt{d}}{||x||_{2}}\right).$ (16)
@@
Figures 9 and 10 show the loss values and gradient norms of the Vanilla configuration in Section 4 and the one using RMSNorms instead of LNs (“RMSNorm” in figures) with 350M parameters. We trained them with $\mathrm{lr}=5.0\times 10^{-4}$ as in Section 4888We tried to train them with $\mathrm{lr}=1.0\times 10^{-3}$ but RMSNorm exploded.. As shown in these figures, RMSNorm faced loss and gradient spikes in a similar location to the ones of Vanilla. These empirical results also indicate that the RMSNorms have the same issue as LNs regarding the exploding gradients.
@@ 6 Conclusion
This paper investigated why LLMs sometimes face loss spikes during their pre-training. Specifically, we focused on gradients of internal layers. We introduced two causes of the exploding gradients during LLM pre-training through theoretical analyses: the shortcut and LNs. We also revealed requirements to prevent the exploding gradients. To satisfy the requirements, we also proposed a method combining simple modifications to embeddings and scaled initialization, a widely used initialization method in LLM pre-training. Experiments indicated that the combination avoided loss spikes. Moreover, we demonstrated that the combined method can train with a relatively larger learning rate, and as a result, it can achieve better performance. We hope our theoretical analyses and proposed method avoid wasting valuable time and computational resources during the LLM constructions.
@@ Appendix B Related Work
To stabilize trainings of Transformer-based neural language models, there have been various discussions on the architecture (Xiong et al., 2020; Liu et al., 2020; Takase et al., 2023; Zeng et al., 2023; Zhai et al., 2023), initialization method (Nguyen & Salazar, 2019; Zhang et al., 2019b; Huang et al., 2020; Wang et al., 2022), training strategy (Zhang et al., 2022; Li et al., 2022), and loss function (Chowdhery et al., 2022; Wortsman et al., 2023).
Xiong et al. (2020) theoretically analyzed gradient scales of each part in Transformers, and indicated that the Pre-LN Transformer is more stable than the Post-LN Transformer, that is the original Transformer architecture (Vaswani et al., 2017). Since the Pre-LN Transformer is more stable than the Post-LN Transformer theoretically and empirically, recent studies mainly have used the Pre-LN Transformer to construct an LLM. We also assume using the Pre-LN Transformer in the analysis on the training dynamics in this paper.
@@
For the initialization methods, Nguyen & Salazar (2019) proposed a strategy to initialize parameters of Transformers with small values to stabilize their training. Zhang et al. (2019b) and Huang et al. (2020) indicated that we can remove layer normalizations in Transformers if we use their proposed initialization methods. Wang et al. (2022) adjusted initial parameter scales based on the number of layers to stabilize the Post-LN Transformer. In this study, we indicated that an initialization method affects the LLM pre-training dynamics. In particular, we indicated that the widely used initialization method (Shoeybi et al., 2020) prevents the explosion due to the shortcut but causes the explosion due to layer normalizations. Moreover, we proved that we can prevent the explosion due to layer normalizations by making the variance of embeddings close to $1$.
@@
Rajbhandari et al. (2020) proposed ZeRO that reduces memory redundancies during the multi GPU training without increasing communication volume. Dao et al. (2022) focused on GPU memory reads/writes, and proposed FlashAttention that accelerates the speed of attention mechanisms in Transformers. To reduce the number of computations in the attention mechanism, Shazeer (2019) proposed the multi-query attention that shares one key and value across all of the attention heads in each layer. Takase & Kiyono (2023) explored several parameter sharing strategies, and indicated that parameter sharing across some layers can achieve comparable performance to the vanilla model with a small number of parameters. Moreover, several studies have explored a better construction way with a limited budget (Izsak et al., 2021; Takase & Kiyono, 2021). We believe that we can take advantage of their findings to make our LLMs more efficient.
@@ Appendix D Comparison with Post-LN Transformer
As described in Section 2.1, recent studies use the Pre-LN Transformer architecture to construct their LLMs because the architecture is more stable. In contrast, some recent studies reported that the Post-LN Transformer, which is the original architecture, can achieve better performance than the Pre-LN if we address the instability issue in the Post-LN, i.e., the vanishing gradient problem (Liu et al., 2020; Takase et al., 2023; Wang et al., 2022). We discuss whether the Pre-LN Transformer entirely underperforms the Post-LN. We conducted experiments on machine translation experiments because previous studies mainly focused on them.
We followed the experimental settings in Takase et al. (2023). Table 4 shows the details of hyper-parameters. We used the WMT English-to-German training dataset (Vaswani et al., 2017; Ott et al., 2018), and evaluated each model in newstest2010-2016. We used the encoder-decoder architecture proposed by Peitz et al. (2019). To stabilize the Post-LN Transformer, we applied DeepNet (Wang et al., 2022) and B2T connection (Takase et al., 2023). We compared them to Scaled Embed, that is, the Pre-LN Transformer with the stabilizing techniques described in this paper.
Table 5 shows the averaged BLEU scores among newstest2010-2016. For the BLEU score calculation, we used SacreBLEU (Post, 2018) to obtain compatible scores (Marie et al., 2021). The signature of SacreBLEU is BLEU+case.mixed+numrefs.1+smooth.exp+tok.13a+version.1.5.0. As shown in this table, we used two learning rates: $\mathrm{lr}=1.0\times 10^{-3}$ and $3.0\times 10^{-3}$. For $\mathrm{lr}=1.0\times 10^{-3}$, DeepNet and B2T connection outperformed Scaled Embed. Thus, the Post-LN Transformer-based methods achieved better performance than the Pre-LN Transformer-based method. This result corresponds to reports in previous studies (Liu et al., 2020; Wang et al., 2022; Takase et al., 2023).
On the other hand, for $\mathrm{lr}=3.0\times 10^{-3}$, Scaled Embed achieved better performance than the others with $\mathrm{lr}=1.0\times 10^{-3}$, and the training of the others failed due to the exploding gradients. This result indicates that the Pre-LN Transformer-based method can achieve better performance if we use a large learning rate. Therefore, the Pre-LN Transformer (with the stabilizing techniques) is more stable than the Post-LN Transformer-based method, and thus, it can achieve better performance when we use a large learning rate that is too large to train the Post-LN Transformers.
@@ Appendix E Pre-training of the Model with 13B Parameters
We conducted pre-trainings of models with 13B parameters to indicate that our modification can stabilize a model with many more parameters than the ones discussed in Section 4. To make this experiment close to a realistic situation, as shown in Table 3, we increased the batch size and the number of updates, and decreased the Adam $\beta_{2}$. In particular, for Adam $\beta_{2}$, most studies have used $0.95$ to stabilize their pre-trainings (Brown et al., 2020; Zhang et al., 2022; Zeng et al., 2023; Biderman et al., 2023; Touvron et al., 2023), and thus, we also used $0.95$ in this experiment. We tried two learning rates: $3.0\times 10^{-4}$, which is the same value in Touvron et al. (2023), and $1.0\times 10^{-4}$.
Figure 11 shows the loss values of each configuration in validation data. As shown in (a) of this figure, the loss value of Vanilla rose from approximately 10000 steps in $\mathrm{lr}=3.0\times 10^{-4}$. Then, the gradient of this model became too large to continue its pre-training. In contrast, the loss value of Scaled Embed consistently decreased. This result indicates that Scaled Embed stabilized the pre-training. We emphasize that the pre-training of Vanilla is essentially unstable even if we use the widely used Adam $\beta_{2}$ value, $0.95$, which is known as the technique to stabilize the pre-training, and our modification is also effective for the stabilization in this realistic situation.
Table 6 shows the perplexities of each configuration in evaluation data. This table indicates that we can achieve better performance when we use a larger learning rate in the same as in Section 5.1. In addition, the perplexities of Scaled Embed were comparable to ones of Vanilla when we used the small learning rate: $\mathrm{lr}=1.0\times 10^{-4}$. These results imply that our modification has no considerable risk in pre-training. Thus, we recommend making the standard deviation of each embedding close to $1$ to stabilize the pre-trainings of LLMs.